# Content Bias in Deep Learning Image Age Approximation: A new Approach Towards better Explainability

### v2

@@ 4 Conclusion
Content bias is an important problem in the context of deep learning image age approximation. Based on the experiments conducted, we have seen that it is difficult for a ‘standard’ CNN to exploit the subtle age traces, even if it is not distracted by content bias in the training data. In this work, we proposed a method that evaluates the influence of image content (based on different types of average images). The functionality of this method was validated using models trained on synthetic images (where content bias can be ruled out). Based on this approach, we showed that ‘standard’ CNNs (AlexNet and SRNet) trained on regular images are strongly dependent on image content. Even if preprocessing is applied to increase the signal-to-noise ratio (i.e., SRNet-res, SRNet-cs and ZhuNet), the image content still plays an important role. Based on the results and in comparison to all variants examined, the image content is least involved in inference with the SRNet-cs. Since it is important that inference is based on comprehensible evidence, we recommend classical approaches to approximating the image age (i.e., [2, 3]) at current stage.

### v3

@@ 4 Conclusion
Content bias is an important problem in the context of deep learning image age approximation. Based on the experiments conducted, we have seen that it is difficult for a ‘standard’ CNN to exploit the subtle age traces, even if it is not distracted by content bias in the training data. In this work, we proposed a method that evaluates the influence of image content (based on different types of average images). The functionality of this method was validated using models trained on synthetic images (where content bias can be ruled out). Based on this approach, we showed that ‘standard’ CNNs (AlexNet and SRNet) trained on regular images are strongly dependent on image content. Image content still plays an important role, even when pre-processing is applied to increase the signal-to-noise ratio (i.e., SRNet-res, SRNet-cs and ZhuNet). Based on the results, image content is least involved in inference with the SRNet-cs. However, since image content is still involved, none of the investigated variants is considered suitable for image age approximation. Thus, we recommend classical image age approximation methods [2, 3] at current stage.